OpenStudy (anonymous):

please help, find four consecutive integers whose sum is -2

OpenStudy (chris):

You can actually set this up as an equation, it's just a bit tricky. So what you know about four consecutive integers is that each integer is one more than the next.

OpenStudy (chris):

So (x) + (x+1) + (x+2) + (x+3) = -2

OpenStudy (chris):

Can you solve it from there?

OpenStudy (chris):

x would be the first integer, and then you know the rest.

OpenStudy (anonymous):

im a little confuse so u have to add all the x together right?

OpenStudy (chris):

Yes, when the problem says to find four consecutive integers that add to get -2, that's what they mean. You know because they are "consecutive", that given a firs integer "x", that the rest follow that pattern - each next integer is 1 greater than the last.

OpenStudy (chris):

(x) + (x+1) + (x+2) + (x+3) = -2 (adding the consecutive integers gets -2) 4x + 6 = -2

OpenStudy (chris):

So what do you get for x?

OpenStudy (chris):

(the first of the four consecutive integers)

OpenStudy (anonymous):

would it be 4x=-8 then u have to divide 8 with 4 right?

OpenStudy (chris):

correct, well, divide -8 by 4 =)